OpenStudy (anonymous):

PLEASE HELP Difference of Even Powers of Two Numbers 64 - y^6

OpenStudy (anonymous):

I get up to 8^2 - (y^3)^2 = (8 + y^3)(8 - y^3)

OpenStudy (anonymous):

then I get lost

OpenStudy (zarkon):

use the formulas for the sum and difference of cubes

OpenStudy (turingtest):

\[x^3+y^3=(x+y)(x^2-xy+y^2)\]\[x^3-y^3=(x-y)(x^2+xy+y^2)\]

OpenStudy (anonymous):

Would I Get (2 + y)(4 - 2y + y^2)(2 - y)(4 + 2y + y^2)?

OpenStudy (turingtest):

look good to me :)

OpenStudy (anonymous):

Tyvm :D

OpenStudy (anonymous):

Love the picture btw

OpenStudy (turingtest):

thanks! your welcome btw